OpenStudy (anonymous):

How do I find a partial derivative of something like sin(4xy)?

OpenStudy (anonymous):

Derivative with respect to what?

OpenStudy (anonymous):

pretend one variable is a constant.

OpenStudy (anonymous):

both x and y

OpenStudy (anonymous):

so for example if you want partial wrt x treat y as a constant

OpenStudy (anonymous):

you would get \[\cos(4xy)\times 4y\]

OpenStudy (anonymous):

Oh, partials: \[\frac{\delta f}{\delta x} = 4ycos(4xy)\] \[\frac{\delta f}{\delta y} = 4xcos(4xy)\]

OpenStudy (anonymous):

Thanks, I didnt know if the chain rule was still in effect.

OpenStudy (anonymous):

Certainly is.

OpenStudy (anonymous):

You treat the variables you aren't taking the partial with respect to as constant applying all the chain rule stuff, etc.